# On the Unknowable Limits to Prediction

### v2

@@
$\quad+\underbrace{[f^{*}(\mathbf{x}_{\text{true}})-f(\mathbf{x}_{\text{true}}\mid y_{\text{true}})]}_{\begin{subarray}{c}\text{Model approximation gain}\\ \text{(Epistemic)}\end{subarray}}$
@@ 2. Observed Target and Features
  • $\delta_{y}$ is measurement error in $y$, with $\mathbb{E}[\delta_{y}]=\mu_{\delta_{y}}$ and $\operatorname{Var}(\varepsilon)=\sigma_{\delta_{y}}^{2}$.
  • $\delta_{\mathbf{x}}$ is measurement error in $\mathbf{x}$, with $\mathbb{E}[\delta_{\mathbf{x}}]=\mu_{\delta_{\mathbf{x}}}$ and variance $\operatorname{Var}(\varepsilon)=\sigma_{\delta_{\mathbf{x}}}^{2}$. $\mathbf{x}_{\text{observed}}$ is a vector with $1\times K$ dimensions, with each $x_{k}\in\mathbf{X}$ for k=$\{1,2,...,K\}$ as each one of the predictors. Note: $\mathbf{x}_{\text{observed}}$ can have different dimensions to $\mathbf{x}_{\text{true}}$.
@@ 3. Predictive Model
where $\delta_{f}$ is defined based on the optimal $\textbf{x}_{\text{true}}$ and $y_{\text{true}}$, and $\mathbb{E}[\delta_{f}]=\mu_{f},\operatorname{Var}[\delta_{f}]=\sigma_{\delta_{f}}$. The difference between $f(\textbf{x}_{\text{true}}\mid y_{\text{true}})$ and $y_{\text{pred}}$ attributes to improvement in $\delta_{\textbf{x}}$ and $\delta_{y}$; further details can be found in Equation LABEL:eq:y_pred_decomposition_refined.
@@
Following these aforementioned definitions, we can decompose the output of a predictive machine which predicts $y_{\text{true}}$ as perfectly as possible, and in the process link it to existent predictions generated by $f()$ to create $y_{pred}$ as follows:
@@
[…]
@@
$Error$ $=\underbrace{\left[{f}(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{observed}}-\delta_{\textbf{x}}\mid y_{\text{observed}})\right]}_{\shortstack{\scriptsize\text{Error due to measurement error in $X$} \\ \\ \scriptsize\text{(Epistemic)}}}$
$\quad+\underbrace{\left[{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{true}}\mid(y_{\text{observed}}-\delta_{y}))\right]}_{\shortstack{\scriptsize\text{Error due to measurement error in $y$ affecting model estimation}\\ \scriptsize\text{(Epistemic)}}}$
@@
$=\underbrace{\left[{f}(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{observed}}-\delta_{\textbf{x}}\mid y_{\text{observed}})\right]}_{\shortstack{\scriptsize\text{Error due to measurement error in $X$} \\ \\ \scriptsize\text{(Epistemic)}}}$
$\quad+\underbrace{\left[{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{true}}\mid(y_{\text{observed}}-\delta_{y}))\right]}_{\shortstack{\scriptsize\text{Error due to measurement error in $y$ affecting model estimation}\\ \scriptsize\text{(Epistemic)}}}$
@@
$=\delta^{\text{error}}_{\textbf{x}}+\delta^{\text{error}}_{y}+\delta_{f}+\varepsilon,$ (S10)
@@
  • $\delta^{\text{error}}_{y}\coloneqq{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{true}}\mid(y_{\text{observed}}-\delta_{y}))$; the difference in predicted values owing to $\delta_{y}$. $\mathbb{E}[\delta^{\text{error}}_{y}]=\mu^{\text{error}}_{y}$ and variance $\operatorname{Var}(\delta^{\text{error}}_{y})=\sigma^{2}_{\delta^{\text{error}}_{y}}$. Similarly, the calculation of $\operatorname{Var}(\delta^{\text{error}}_{y})$ depends on the specific form of $f$, the distribution of $y_{\text{observed}}\text{ and }y_{\text{observed}}-\delta_{y}$, and their dependencies.
@@
$=\mu^{\text{error}}_{\textbf{x}}+\mu^{\text{error}}_{y}+\mu_{\delta_{f}}.$ (S11)

### v3

@@
$=\underbrace{[f^{*}(\mathbf{x}_{\text{true}})-f(\mathbf{x}_{\text{true}}\mid y_{\text{true}})]}_{\begin{subarray}{c}\text{Model approximation gain}\\ \text{(Epistemic)}\end{subarray}}$
@@ 2. Observed Target and Features
  • $\delta_{y}$ is measurement error in $y$, with $\mathbb{E}[\delta_{y}]=\mu_{\delta_{y}}$ and $\operatorname{Var}(\delta_{y})=\sigma_{\delta_{y}}^{2}$.
  • $\delta_{\mathbf{x}}$ is measurement error in $\mathbf{x}$, with $\mathbb{E}[\delta_{\mathbf{x}}]=\mu_{\delta_{\mathbf{x}}}$ and variance $\operatorname{Var}(\delta_{\mathbf{x}})=\sigma_{\delta_{\mathbf{x}}}^{2}$. $\mathbf{x}_{\text{observed}}$ is a vector with $1\times K$ dimensions, with each $x_{k}\in\mathbf{X}$ for $k=\{1,2,...,K\}$ as each one of the predictors. Note: $\mathbf{x}_{\text{observed}}$ can have different dimensions to $\mathbf{x}_{\text{true}}$.
@@ 3. Predictive Model
where $\delta_{f}$ is defined based on the optimal $\textbf{x}_{\text{true}}$ and $y_{\text{true}}$, and $\mathbb{E}[\delta_{f}]=\mu_{\delta_{f}},\operatorname{Var}[\delta_{f}]=\sigma_{\delta_{f}}$. The difference between $f(\textbf{x}_{\text{true}}\mid y_{\text{true}})$ and $y_{\text{pred}}$ attributes to improvement in $\delta_{\textbf{x}}$ and $\delta_{y}$; further details can be found in Equation LABEL:eq:y_pred_decomposition_refined.
@@
Following these aforementioned definitions, we can decompose the output of a predictive machine which predicts $y_{\text{true}}$ as perfectly as possible, and in the process link it to existent predictions generated by $f()$ to create $y_{\text{pred}}$ as follows:
@@
$\quad+\underbrace{[f(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-f(\textbf{x}_{\text{true}}+\delta_{\textbf{x}}\mid y_{\text{observed}})]}_{\scriptsize\shortstack{\text{Measurement gain from} $\mathbf{x}$ \\ \text{(Epistemic)}}}$
@@
$Error$ $=\underbrace{\left[{f}(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{observed}}-\delta_{\textbf{x}}\mid y_{\text{observed}})\right]}_{\shortstack{\scriptsize\text{Error due to measurement error in $\mathbf{x}$} \\ \\ \scriptsize\text{(Epistemic)}}}$
$\quad+\underbrace{\left[{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}}-\delta_{y})\right]}_{\shortstack{\scriptsize\text{Error due to measurement error in $y$ affecting model estimation}\\ \scriptsize\text{(Epistemic)}}}$
@@
$=\underbrace{\left[{f}(\textbf{x}_{\text{observed}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{observed}}-\delta_{\textbf{x}}\mid y_{\text{observed}})\right]}_{\shortstack{\scriptsize\text{Error due to measurement error in $\mathbf{x}$} \\ \\ \scriptsize\text{(Epistemic)}}}$
$\quad+\underbrace{\left[{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}}-\delta_{y})\right]}_{\shortstack{\scriptsize\text{Error due to measurement error in $y$ affecting model estimation}\\ \scriptsize\text{(Epistemic)}}}$
@@
$=\delta^{\text{error}}_{\mathbf{x}}+\delta^{\text{error}}_{y}+\delta_{f}+\varepsilon,$ (S10)
@@
  • $\delta^{\text{error}}_{y}\coloneqq{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}})-{f}(\textbf{x}_{\text{true}}\mid y_{\text{observed}}-\delta_{y})$; the difference in predicted values owing to $\delta_{y}$. $\mathbb{E}[\delta^{\text{error}}_{y}]=\mu^{\text{error}}_{y}$ and variance $\operatorname{Var}(\delta^{\text{error}}_{y})=\sigma^{2}_{\delta^{\text{error}}_{y}}$. Similarly, the calculation of $\operatorname{Var}(\delta^{\text{error}}_{y})$ depends on the specific form of $f$, the distribution of $y_{\text{observed}}\text{ and }y_{\text{observed}}-\delta_{y}$, and their dependencies.
@@
$=\mu^{\text{error}}_{\textbf{x}}+\mu^{\text{error}}_{y}+\mu_{f}.$ (S11)